# Pázmány Péter sétány 1/A, 1117 Budapest, Hungary

### v1

@@ 3.4 Evaluation representations
therefore we can lift any $\mathfrak{o}_{4}$ irreps to $Y^{+}(4)$ irreps. Since $\mathfrak{o}_{4}\cong\mathfrak{sl}_{2}\oplus\mathfrak{sl}_{2}$, any 2-tuple $(s_{L},s_{R})$ defines an $\mathfrak{o}_{4}$ highest weight irrep ($s_{L/R}$ are the spins of the $\mathfrak{sl}_{2}$ irreps). Let us denote $V(s_{L},s_{R})$ the corresponding $Y^{+}(4)$ irrep.
Let us parametrize the $\mathfrak{o}_{4}=\mathfrak{sl}_{2}\oplus\mathfrak{sl}_{2}$ generators as
@@
Using the relation (3.18), the explicit form of the $K$-matrix can easily be written down. We have listed it together with related quantities relevant for the present section in appendix A.
Figure 4: The graphical presentation of the K-matrices, boundary states. The red lines correspond to the auxiliary space ($i,j=1,\dots 4$). The boundary space denoted by orange lines. The intersection of two red lines denote the R-matrix $R(u)$. Here, we used the shorthand notations $u^{\pm}=u\pm 1/2$ for the spectral parameter dependence. The rounded rectangles are the two-site boundary operators $\psi^{(0)}_{i,j}(u)=\mathbf{K}_{5-j,i}(u)$ for the defining representation of $\mathfrak{gl}_{4}$
The corresponding boundary state can be obtained using the fusion procedure (the sketch of the fusion is shown in figure 4). If the quantum space were the defining representation then the boundary state would be constructed from the two-site operators $\psi^{(0)}_{i,j}(u)=\mathbf{K}_{5-j,i}(u)$. The fused two-site operator is
@@ 4 Dressing
We can apply the dressing procedure to the one-dimensional representation i.e. when the $K$-matrix is
@@
where $\varphi_{i,j}$ was given in eqn. (A.5). After the dressing the representation is $\mathcal{V}_{B}=L(\lambda|\xi)$ with the following dressed $K$-matrix and associated integrable boundary state
@@
4.1 Generalized dressing formulas for MPS’s
In the previous section we saw that for special MPS’s obtained by dressing the K-matrix corresponding to the one-dimensional representation of $Y^{+}(4)$ by $L(\lambda|\xi)$ there exists a simple, closed formula for the overlap with Bethe eigenstates, because the MPS can be expressed as an action of a transfer matrix on the simple two site product state $|\delta\rangle$.
@@ 4.1 Generalized dressing formulas for MPS’s
Our goal is to find the generalized dressing formulas and the overlap ratios for the states $\langle\mbox{MPS}_{2s+1}^{L/R}|$ which will allow us to fill the last gap in the proof of the expression for the tree level one-point functions of scalars in the D3-D5 domain wall version of ${\cal N}=4$ SYM. We follow the strategy of DeLeeuw:2019ohp where the overlap ratio was determined for simpler MPS’s describing the tree-level one-point functions in the $SU(3)$ sub-sector. In the first subsection we review the method of DeLeeuw:2019ohp , next we apply it to the $\langle\mbox{MPS}_{2s+1}^{L/R}|$.
@@ 4.2 Methodology
In the following we describe our methodology with a general example. Let us pick a $Y^{+}(4)$ irrep $V\equiv V^{(1)}$ with the corresponding $K$-matrix $\mathbf{K}_{i,j}(u)\equiv\mathbf{K}_{i,j}^{(1)}(u)$ and boundary state $\langle\Psi|\equiv\langle\Psi^{(1)}|$ which satisfy the KT-relation, The matrix product state $\langle\mbox{MPS}|\equiv\langle\mbox{MPS}^{(1)}|$ is defined from the boundary state by taking the twisted trace in the boundary space.
@@ 4.3 Generalized dressing for fixed spin
For half integer spins we cannot follow the previous strategy since there is no finite dimensional $\mathfrak{gl}_{4}$ irrep for which the ratio of highest weights (4.16) is the same as (3.29) for half integer spins. This suggests that the MPS’s for half integer spins cannot be expressed through a two-site state. Therefore, in this case we look for another type of dressing formula. Our goal will be to relate the general half integer spin states $\langle\mbox{MPS}_{2s+1}^{R}|$ to the 1/2 spin states $\langle\sigma^{R}|\equiv\langle\mbox{MPS}_{2}^{R}|$ and $\langle\sigma^{L}|\equiv\langle\mbox{MPS}_{2}^{L}|$.
@@
4.4 Universality of the generalized dressing formulas
Since our calculation is based on the $K$-matrix (which is universal, i.e., do not depend on the representation of the quantum space) the generalized dressing formulas (4.32) and (4.39) are true for any representations of the quantum space (up to the normalization scalars). To demonstrate this universality let us pick another example. Let us change the quantum space to defining representation of $\mathfrak{gl}_{4}$ (usual $SU(4)$ spin chain). Since the algebra is the same (RTT and the KT is the same), we only need to modify the explicit form the Lax-operators
@@ 4.4 Universality of the generalized dressing formulas
In summary, we learn that the generalized dressing formula (4.32) and (4.39) are universal, i.e. they are true for any representations, any inhomogeneities and any twists. This implies that, when searching for overlap formulas (even with brute force numerical methods), it is always advisable to use the simplest possible quantum space, since the formula is universal. For the given quantum space the formula is modified only by the normalization pre-factors which are easy to fit at the end.
5 Application to the D3-D5 domain wall
@@
where $\mathcal{R}\in GL(4)$ and $\Delta(\mathcal{R})$, $\mathcal{R}^{\lambda}$ are its representation in the quantum and auxiliary spaces, more precisely
@@
In the untwisted case the transfer matrices also has $GL(4)$ symmetry:
@@
This symmetry can be used on (4.32) to get generalized dressing formula for the ”rotated” MPS’s
@@
where we introduced the rotated MPS and two-site states
$\langle\mbox{MPS}^{\mathcal{R}}|=\langle\mbox{MPS}|\Delta(\mathcal{R}),\quad\langle\delta^{\mathcal{R}}|=\langle\delta|\Delta(\mathcal{R}).$ (5.6)
The rotation for the other type of generalized dressing (4.39) is analogous. If the original MPS and two-site state are given by the formulas (3.23) and (4.12) then the rotated ones are
@@
therefore the overlap of the state $\langle\mbox{MPS}^{L,\mathcal{R}}|$ can be expressed with the overlap of the other state $\langle MPS^{R,\mathcal{R}}|$ as
@@
Notice that the calculations of this section cannot be applied in the twisted case, since the twisted transfer matrices do not have $GL(4)$ symmetry and the Bethe roots do not have pair structure, i.e., $Q_{k}(z)\neq Q_{k}(-z)$. Due to the lack of the $GL(4)$ symmetry, it is not possible to use “rotations” to relate the overlaps $\langle\mbox{MPS}_{2s+1}^{L}|\bar{u}\rangle$ to the overlaps $\langle\mbox{MPS}_{2s+1}^{R}|\bar{u}\rangle$, i.e., the dressing formula only gives the sum of the overlaps $\langle\mbox{MPS}_{2s+1}^{L}|\bar{u}\rangle+\langle\mbox{MPS}_{2s+1}^{R}|\bar{u}\rangle$. Due to the lack of pair structure, the linear combination $R_{s}$ does not simplify. This might improve if we introduce the full Q-system Tsuboi:1998ne , for reviews see Gromov:2019icz ; Kazakov:2018ugh ; Levkovich-Maslyuk:2019awk $\mathcal{Q}_{\emptyset},\mathcal{Q}_{i},\mathcal{Q}_{ij},\mathcal{Q}_{ijk},\mathcal{Q}_{1234}$ (our Q’s are $Q_{1}=\mathcal{Q}_{1}$, $Q_{2}=\mathcal{Q}_{12}$, $Q_{3}=\mathcal{Q}_{123}$). We stress again that the formula (4.32) is valid even for the twisted case.
@@ 6 Discussion and Conclusion
Provided one can determine the overlap between the Bethe eigenstates and the MPS corresponding to the trivial representation of $Y^{+}(4)$, and provided that the left and the right version of the general MPS can be related, the dressing formulas give access to a closed overlap formula for any MPS of evaluation representation type. In the case where the twist vanishes the former overlap can be read off from Gombor:2021hmj ; Gombor:2023bez where the overlap was determined for any two site product state for any GL(N) spin chain. Devising a method which allows one to find the same overlap in the presence of twists constitutes an interesting open problem which could possibly be approached by means of the functional SoV approach Sklyanin:1995bm ; Gromov:2016itr ; Ekhammar:2023iph . In the case where twists are present, however, there is no simple relation between the left and the right version of the MPS. Yet our dressing formulas give rise to a non-trivial summation rule for right and left MPS overlaps. Furthermore the overlaps with the trivial representation in the twisted case may simplify if one introduces the full $Q$-system Tsuboi:1998ne ; Gromov:2019icz ; Kazakov:2018ugh ; Levkovich-Maslyuk:2019awk and not only the limited set of $Q$-functions used here.
One may wonder if the general evaluation representations $V(s_{R},s_{L})$ and their corresponding integrable matrix product states have an interpretation within the AdS/CFT correspondence. An obvious place to search for such an interpretation would be the non-supersymmetric D3-D7 domain wall version of ${\cal N}=4$ SYM, which is characterized by the scalar fields attaining vevs which are given by two different representations of $\mathfrak{su}(2)$ Kristjansen:2012tn . A glance at the eqns (5.11) immediately shows that this interpretation can not be correct as the mapping of the spin chain states to a set of three complex scalar fields and their complex conjugates (or rather six real scalar fields) only works for the representations $V_{R}(s)$ and $V_{L}(s)$. This is in accordance with the fact that the former D3-D7 domain wall version of ${\cal N}=4$ SYM was shown to be non integrable deLeeuw:2019sew .
@@
It would likewise be interesting to extend the results of the present paper to other spin chains and in particular other reflection algebras. The extension of overlap formulas to other reflection algebras, albeit in the case of vanishing twists, is addressed in the paper Gombor:2024Oct . Whereas each reflection algebra has to be treated independently the resulting formulas are highly universal.
@@ Acknowledgements.
C.K. was supported in part by DFF-FNU through grant number 1026-00103B. V.M. was supported by the STFC under the grant ST/X508809/1. X.Q. was supported by CSC through grant number 202207940019, he also thanks Yau center of Southeast University for hospitality. T.G. was supported by the NKFIH grant PD142929 and the János Bolyai Research Scholarship of the Hungarian Academy of Science.

### v2

@@ 3.5 Evaluation representations
therefore we can lift any $\mathfrak{o}_{4}$ irreps to $Y^{+}(4)$ irreps. Since $\mathfrak{o}_{4}\cong\mathfrak{sl}_{2}\oplus\mathfrak{sl}_{2}$, any 2-tuple $(s_{L},s_{R})$ defines an $\mathfrak{o}_{4}$ highest weight irrep ($s_{L/R}$ are the spins of the $\mathfrak{sl}_{2}$ irreps). Let us denote by $V(s_{L},s_{R})$ the corresponding $Y^{+}(4)$ irrep.
Furthermore, let us parametrize the $\mathfrak{o}_{4}=\mathfrak{sl}_{2}\oplus\mathfrak{sl}_{2}$ generators as
@@
Using the relation (3.25), the explicit form of the $K$-matrix can easily be written down. We have listed it together with related quantities of relevance for the present section in appendix A.
Figure 4: The graphical presentation fusion leading to the boundary state. The red lines correspond to the auxiliary space ($i,j=1,\dots 4$). The boundary space is denoted by orange lines. The intersection of two red lines denote the R-matrix $R(u)$. Here, we used the shorthand notations $u^{\pm}=u\pm 1/2$ for the spectral parameter dependence. The rounded rectangles are the two-site boundary operators $\psi^{(0)}_{i,j}(u)=\mathbf{K}_{5-j,i}(u)$ for the defining representation of $\mathfrak{gl}_{4}$
The corresponding boundary state can be obtained using the fusion procedure sketched in figure 4. If the quantum space were the defining representation the boundary state would be constructed from the two-site operators $\psi^{(0)}_{i,j}(u)=\mathbf{K}_{5-j,i}(u)$. The fused two-site operator is
@@ 4 Dressing
We can apply the dressing procedure to the one-dimensional representation, i.e. when the $K$-matrix is
@@
where $\varphi_{i,j}$ is given in eqn. (A.5). After the dressing the representation is $\mathcal{V}_{B}=L(\lambda|\xi)$ with the following dressed $K$-matrix and associated integrable boundary state
@@
4.1 Generalized dressing formulas for MPSs
In the previous section we saw that for special MPSs obtained by dressing the K-matrix corresponding to the one-dimensional representation of $Y^{+}(4)$ by $L(\lambda|\xi)$ there exists a simple, closed formula for the overlap with Bethe eigenstates, because the MPS can be expressed as an action of a transfer matrix on the simple two site product state $|\delta\rangle$.
@@ 4.1 Generalized dressing formulas for MPSs
Our goal is to find the generalized dressing formulas and the overlap ratios for the states $\langle\mbox{MPS}_{2s+1}^{L/R}|$ which will allow us to fill the last gap in the proof of the expression for the tree level one-point functions of scalars in the D3-D5 domain wall version of ${\cal N}=4$ SYM. We follow the strategy of DeLeeuw:2019ohp where the overlap ratio was determined for simpler MPSs describing the tree-level one-point functions in the $SU(3)$ sub-sector. In the first subsection we review the method of DeLeeuw:2019ohp , next we apply it to the $\langle\mbox{MPS}_{2s+1}^{L/R}|$.
@@ 4.2 Methodology
In the following we describe our methodology with a general example. Let us pick a $Y^{+}(4)$ irrep $V\equiv V^{(1)}$ with the corresponding $K$-matrix $\mathbf{K}_{i,j}(u)\equiv\mathbf{K}_{i,j}^{(1)}(u)$ and boundary state $\langle\Psi|\equiv\langle\Psi^{(1)}|$ which satisfy the KT-relation. The matrix product state $\langle\mbox{MPS}|\equiv\langle\mbox{MPS}^{(1)}|$ is defined from the boundary state by taking the twisted trace in the boundary space.
@@ 4.3 Generalized dressing for fixed spin
For half integer spins we cannot follow the previous strategy since there is no finite dimensional $\mathfrak{gl}_{4}$ irrep for which the ratio of highest weights (4.16) is the same as (3.36) for half integer spins. This suggests that the MPSs for half integer spins cannot be expressed through a two-site state. Therefore, in this case we look for another type of dressing formula. Our goal will be to relate the general half integer spin states $\langle\mbox{MPS}_{2s+1}^{R}|$ to the 1/2 spin states $\langle\sigma^{R}|\equiv\langle\mbox{MPS}_{2}^{R}|$ and $\langle\sigma^{L}|\equiv\langle\mbox{MPS}_{2}^{L}|$.
@@
4.4 Universality of the generalized dressing formulas wrt. the quantum space
Since our calculation is based on the $K$-matrix (which is universal, i.e., do not depend on the representation of the quantum space) the generalized dressing formulas (4.32) and (4.39) are true for any representations of the quantum space (up to the normalization scalars). To demonstrate this universality let us pick another example. Let us change the quantum space to defining representation of $\mathfrak{gl}_{4}$ (usual $SU(4)$ spin chain). Since the algebra is the same (RTT and the KT are the same), we only need to modify the explicit form the Lax-operators
@@
In summary, we learn that the generalized dressing formula (4.32) and (4.39) are universal, i.e. they are true for any representations. This implies that, when searching for overlap formulas (even with brute force numerical methods), it is always advisable to use the simplest possible quantum space, since the formula is universal. For the given quantum space the formula is modified only by the normalization pre-factors which are easy to fit at the end.
4.5 Universality of the generalized dressing formulas wrt. the twist
Until now we focused on diagonal twists but in this section we will show that generalized dressing formulas can be derived for any twist.
@@
where $\mathcal{R}\in GL(4)$, and $\Delta(\mathcal{R})$ and $\mathcal{R}^{\lambda}$ are its representations in the quantum and auxiliary spaces, respectively. More precisely
@@
From the symmetry equation of the monodromy matrix, eqn. (4.46) we obtain
$\mathrm{Tr}_{0}\left(T_{0}(u)G_{0}\right)\Delta(\mathcal{R})=\Delta(\mathcal{R})\mathrm{Tr}_{0}\left(T_{0}(u)\mathcal{R}_{0}^{-1}G_{0}\mathcal{R}_{0}\right),$ (4.49)
Therefore,
$\mathcal{T}^{G}(u)\Delta(\mathcal{R})=\Delta(\mathcal{R})\mathcal{T}^{G_{\mathcal{R}}}(u),$ (4.50)
where $G_{\mathcal{R}}=\mathcal{R}^{-1}G\mathcal{R}$ and the superscript denotes the twist matrix. This relation can be also generalized to the other transfer matrices
$\widehat{\mathcal{T}}^{G}(u)\Delta(\mathcal{R})=\Delta(\mathcal{R})\widehat{\mathcal{T}}^{G_{\mathcal{R}}}(u),\quad\mathcal{T}_{\lambda}^{G}(u)\Delta(\mathcal{R})=\Delta(\mathcal{R})\mathcal{T}_{\lambda}^{G_{\mathcal{R}}}(u),$ (4.51)
where
$\widehat{\mathcal{T}}^{G}(u)=\mathrm{Tr}_{0}\left(\widehat{T}_{0}(u)\widehat{G}_{0}\right),\quad\mathcal{T}_{\lambda}^{G}(u)=\mathrm{Tr}_{0}\left(T_{0}^{\lambda}(u)G_{0}^{\lambda}\right)$ (4.52)
Let us introduce a rotated MPS by
$\langle\mbox{MPS}^{\mathcal{R}}|=\langle\mbox{MPS}|\Delta(\mathcal{R}),$ (4.53)
If an MPS is compatible with a certain twist matrix i.e.
$\langle\mathrm{MPS}|\mathcal{T}^{G}(u)=\langle\mathrm{MPS}|\widehat{\mathcal{T}}^{G}(-u),$ (4.54)
then the rotated MPS is compatible with the rotated twist
$\langle\mathrm{MPS}^{\mathcal{R}}|\mathcal{T}^{G_{\mathcal{R}}}(u)=\langle\mathrm{MPS}^{\mathcal{R}}|\widehat{\mathcal{T}}^{G_{\mathcal{R}}}(-u).$ (4.55)
If our dressing formulas are satisfied for a pair of MPS and twist $(\langle\mathrm{MPS}|,G)$ they are also satisfied for the rotated versions $(\langle\mathrm{MPS}^{\mathcal{R}}|,G_{\mathcal{R}})$. This means that our dressing formulas also work for non-diagonal twists. We just need to make sure that we choose the K-matrix and MPS that are compatible with the twist (the equations (3.5) and (3.6) are satisfied).
In practical terms, if we have a general twist and a compatible K-matrix (eq. (3.5)) then we can use a GL(4) transformation like in (4.53) to diagonalize the twist matrix G. In the rotated frame the twist matrix G is diagonal, therefore the dressing formulas are true in that frame. Finally, we can rotate back and get the dressing formulas in the original frame. These formulas will involve transfer matrices with modified twists. Any twist can be obtained from a diagonal one by a GL(4) transformation, therefore generalized dressing formulas exist for any twist.
5 Application to the D3-D5 domain wall
In this section we finally return to the overlap of relevance for the D3-D5 domain wall described in the introduction. Our aim is to derive the overlap $\langle\mbox{MPS}_{k}|\bar{u}\rangle$ between the matrix product state given by eqns. (1.1) and (1.2) and the eigenstates of the integrable $\mathfrak{so}(6)$ spin chain. We first exploit the $GL(4)$ symmetry of the monodromy and transfer matrices to perform a specific $GL(4)$ rotation which transforms the dressing formulas (4.32) and (4.39) into formulas involving exactly the integrable matrix product state of interest to us. Secondly, by another $GL(4)$ rotation we relate the left and the right version of the overlap for the integer spin case. Straightforward manipulations of the dressing formulas subsequently allow us to recover the result previously derive by numerical investigations DeLeeuw:2018cal .
As in the previous section we start from the symmetry relation for the monodromy matrix, eqn. (4.46) with a $GL(4)$ rotation $\mathcal{R}$ represented in the auxiliary and quantum spaces as in (4.47) and (4.48). In the untwisted case the transfer matrices also have $GL(4)$ symmetry:
@@ 5 Application to the D3-D5 domain wall
This symmetry can be applied to (4.32) to get the generalized dressing formula for the ”rotated” MPSs
@@
where the rotated MPSs are defined as in eqn. (4.53) and similarly for the rotated two-site state. The rotation for the other type of generalized dressing formula (4.39) is analogous. If the original MPS and two-site state are given by the formulas (3.30) and (4.12) then the rotated ones are
@@
therefore the overlap of the state $\langle\mbox{MPS}^{L,\mathcal{R}}|$ can be expressed with the overlap of the other state $\langle\mbox{MPS}^{R,\mathcal{R}}|$ as
@@
Notice that the detailed calculations of on-shell overlaps of this section cannot be applied in the twisted case since the twisted transfer matrices do not have $GL(4)$ symmetry (rather the $GL(4)$ rotations transform between different twists), and the selection rules do no imply pairing of Bethe roots as explained in section 3.4. Due to the lack of GL(4) symmetry, it is not possible to use “rotations” to relate the overlaps $\langle\mathrm{MPS}_{2s+1}^{L}|\bar{u}\rangle$ to the overlaps $\langle\mathrm{MPS}_{2s+1}^{R}|\bar{u}\rangle$, i.e. the dressing formula only gives the sum of the overlaps $\langle\mathrm{MPS}_{2s+1}^{L}|\bar{u}\rangle+\langle\mathrm{MPS}_{2s+1}^{R}|\bar{u}\rangle$. Furthermore, due to the lack of pair structure, the linear combination $R_{s}$ does not simplify.
@@ 6 Discussion and Conclusion
Provided one can determine the overlap between the Bethe eigenstates and the MPS corresponding to the trivial representation of $Y^{+}(4)$, and provided that the left and the right version of the general MPS can be related, the dressing formulas give access to a closed overlap formula for any MPS of evaluation representation type. In the case where the twist vanishes the former overlap can be read off from Gombor:2021hmj ; Gombor:2023bez where the overlap was determined for any two site product state for any GL(N) spin chain. Devising a method which allows one to find the same overlap in the presence of twists constitutes an interesting open problem which could possibly be approached by means of the functional SoV approach Sklyanin:1995bm ; Gromov:2016itr ; Ekhammar:2023iph . In the case where twists are present, however, there is no simple relation between the left and the right version of the MPS. Yet our dressing formulas give rise to a non-trivial summation rule for right and left MPS overlaps. Furthermore the overlaps with the trivial representation in the twisted case may simplify if one introduces the full $\mathcal{Q}$-system Tsuboi:1998ne ; Gromov:2019icz ; Kazakov:2018ugh ; Levkovich-Maslyuk:2019awk and not only the limited set of $Q$-functions used here.
One may wonder if the general evaluation representations $V(s_{R},s_{L})$ and their corresponding integrable matrix product states have an interpretation within the AdS/CFT correspondence. An obvious place to search for such an interpretation would be the non-supersymmetric D3-D7 domain wall version of ${\cal N}=4$ SYM, which is characterized by the scalar fields attaining vevs which are given by two different representations of $\mathfrak{su}(2)$ Kristjansen:2012tn . A glance at the eqn. (5.7) immediately shows that this interpretation can not be correct as the mapping of the spin chain states to a set of three complex scalar fields and their complex conjugates (or rather six real scalar fields) only works for the representations $V_{R}(s)$ and $V_{L}(s)$. This is in accordance with the fact .tethat the former D3-D7 domain wall version of ${\cal N}=4$ SYM was shown to be non integrable deLeeuw:2019sew .
@@
It would likewise be interesting to extend the results of the present paper to other spin chains and in particular other reflection algebras. The extension of overlap formulas to other reflection algebras, albeit in the case of vanishing twists, is addressed in the paper Gombor:2024iix . Whereas each reflection algebra has to be treated independently the resulting formulas are highly universal.
@@ Acknowledgements.
T.G. was supported by the NKFIH grant PD142929 and the János Bolyai Research Scholarship of the Hungarian Academy of Science. C.K. was supported in part by DFF-FNU through grant number 1026-00103B. V.M. was supported by the STFC under the grant ST/X508809/1. X.Q. was supported by CSC through grant number 202207940019, he also thanks Yau center of Southeast University for hospitality.